OpenStudy (anonymous):

how do you simplify sin^4\theta+sin^2\theta cos^2\theta

OpenStudy (anonymous):

let sin/theta = s and cos/theta = c for the sake of these equations\[s^4 + s^2c^2\]factorising s^2 we get\[s^2(s^2+c^2)\]this is the identity\[\cos ^2 + \sin^2 =1\]

OpenStudy (anonymous):

Therefore answer is simply\[\sin^2\]

mathslover (mathslover):

Good work @hea , great to see a new user helping someone. Well done, @Sophiel44 so the answer will be : \(\sin^2 \theta\) , have you got the method hea suggested?

OpenStudy (anonymous):

yes I have thanks!

mathslover (mathslover):

Also, \(\color{blue}{\mathsf{Welcome}} \space \color{orange}{\cal{To}} \space \color{red}{\mathsf{OpenStudy!!!}}\)

OpenStudy (anonymous):

How would you do this one : \[\sin ^{4}\theta+2\sin ^{2}\theta \cos ^{2}\theta +\cos ^{4}\]

mathslover (mathslover):

\(\sin^4 \theta + 2\sin^2 \theta \cos^2 \theta + \cos^4 \theta \) can be written as : \((\sin^2 \theta)^2 + 2\sin^2 \theta \cos^2 \theta + (\cos^2 \theta)^2\) right?

OpenStudy (anonymous):

yes

mathslover (mathslover):

Put : \(\sin^2 \theta = x\) and \(\cos ^2 \theta = y\) We get now : \(x^2 + 2 xy + y^2\)

OpenStudy (anonymous):

then what?

mathslover (mathslover):

Can you simplify \(x^2 +2xy+ y^2\) , I mean factorize it.

OpenStudy (anonymous):

yes, (x+y) (x+y)

mathslover (mathslover):

or (x+y)^2 ?

OpenStudy (anonymous):

oh yes

mathslover (mathslover):

now put x and y as \(\sin^2 \theta\) and \(\cos^2 \theta\) respectively. What you get?

OpenStudy (anonymous):

=1

OpenStudy (anonymous):

I see how this works, thank you very much! :)

mathslover (mathslover):

That's it, Well done sophie. Really great. Best of luck for your journey in OpenStudy ... Keep helping and getting help on OpenStudy .

OpenStudy (anonymous):

officially a fan

mathslover (mathslover):

best of luck again and thanks!